# On quantum learning algorithms for noisy linear problems

### v1

@@ 3.0.3 RLWE to LWE
where ${\bm{a}}_{j}=(\mathsf{M}_{\phi}(a)_{j})^{T}=\big{(}\!\left<{\bm{a}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{a}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{a}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$. We see that the operation is efficiently computable as any power of $\mathbf{P}$ can be classically pre-computed by knowing the fixed polynomial $\phi(X)$.
Now let ${\bm{s}}=\mathsf{V}(s)$ and ${\bm{e}}_{a}=(e_{a,0},\ldots,e_{a,n-1})=\mathsf{V}(e_{a})$. By Lemma 1, Eq. (8), and the linearity of $\mathsf{V}$, we have
$\mathsf{V}(s\cdot a+e_{a}\bmod\phi)$ $=$ $\mathsf{V}(s)\cdot\mathsf{M}_{\phi}(a)+\mathsf{V}(e_{a})$
@@
$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}s(X)\!\cdot\!a(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
$\xmapsto{U_{\phi,j}}$ $\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle\lvert\ldots,\,\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q,\,\ldots\rangle\enspace.$
Ignoring or uncomputing irrelevant information, we have a quantum LWE sample with secret ${\bm{s}}$:
$\lvert{\bm{a}}_{j}\rangle\lvert\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q\rangle\enspace.$
Alternatively, interchanging the role of $a(X)$ and $s(X)$ in the transformation, we have a quantum LWE sample with secret ${\bm{s}}_{j}$:
$\lvert a(X)\rangle\lvert 0\rangle\lvert\big{[}a(X)\!\cdot\!s(X)+e_{a}(X)\bmod\phi\big{]}\bmod q\rangle$
$\xmapsto{\hskip 5.23529pt\mathsf{V}\hskip 5.23529pt}$ $\lvert{\bm{a}}\rangle\lvert{\bm{0}}\rangle\lvert\left<{\bm{a}},{\bm{s}}_{0}\right>+e_{a,0}\bmod q,\ldots,\left<{\bm{a}},{\bm{s}}_{n-1}\right>+e_{a,n-1}\bmod q\rangle$
where ${\bm{s}}_{j}:=(\mathsf{M}_{\phi}({\bm{s}})_{j})^{T}=\big{(}\!\left<{\bm{s}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{s}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{s}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}$.
4 On SIS algorithm
In [12, Appendix B], a quantum sample for SIS is defined by,
$\lvert\Psi^{\textup{sis}}\rangle=\frac{1}{\sqrt{q^{nm}}}\sum_{\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}}\lvert\mathbf{A}\rangle\lvert\mathbf{A}{{\bm{v}}}\bmod q\rangle$ (9)
where $\mathbf{A}=[\,{\bm{a}}_{0}\,|\,\cdots\,|\,{\bm{a}}_{n-1}\,]$, ${\bm{a}}_{i}\in\mathbb{Z}_{q}^{m\times 1}$, and ${\bm{v}}\in\mathbb{Z}^{n}$ is a short secret vector such that $0<\|{\bm{v}}\|\leq\beta$. The authors have developed a polynomial time quantum algorithm for finding ${\bm{v}}$ by using $O(2n\beta)$ quantum samples.
The detailed description of the algorithm is not covered here. Instead, let us focus on the problem itself. From the definition of the SIS problem in Section 2.1.4, we are given an under-determined system of equations ($m<n$). Hardness of the SIS problem comes from the fact that among the exponentially many solutions ($q^{n-m}$), there exists no known way to choose a short one within polynomial time. Having ability to prepare the quantum sample multiple times for the same ${\bm{v}}$ thus already means their is a way to choose the short solution out of many solutions. On the other hand, the hardness of LWE comes from difficulty of solving noisy linear systems. Distinct LWE algorithms may require different number of noisy equations, but accessing more than one noisy equation is not implausible.
Assuming the SIS quantum sample in Eq. (9) is obtainable, a classical algorithm can be designed. By classical, we mean no quantum process is required except measurements. Let $t=O(n/m)$, and for given $t$ samples just measure the state to have
@@ 4 On SIS algorithm
Notice that the total number of linear equations we have is $mt$, which is greater than $n$. With high probability this system of equations has the unique solution ${\bm{v}}$ in which Gaussian elimination can be applied to solve.
5 On the Divide-and-Conquer algorithm
Song et al. have published a paper in 2022 which improves the result of [12] assuming a resized sample [19]. In a nutshell, the improved algorithm and one in Section 2.3 work similarly except the samples. Therefore, here we drop the details on the algorithm and jump right into the discussion on Eq. (9) in [19], which introduces the resized sample. Recall first that an LWE sample is $({\bm{a}},b=\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$, $|e|\leq k$.
5.1 Resized sample assumption
Let $\mathcal{V}_{j}\subseteq\mathbb{Z}_{q}$, ${\bm{s}}=(s_{0},\ldots,s_{n-1})\in\mathbb{Z}_{q}^{n}$, and assume the following sample is obtainable,
$\lvert\psi_{\mathcal{V}_{j}}\rangle=\frac{1}{\sqrt{|\mathcal{V}_{j}|}}\sum_{a_{j}^{\prime}\in\mathcal{V}_{j}}\lvert a_{j}^{\prime}\rangle\lvert a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q\rangle,\quad|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)\ll q$ (10)
where $e_{j}^{\prime}$ can be considered as an error from some distribution originated from the original LWE sample.
Apparently some more assumptions are required (including QRAM) for the above statement to hold, but let us take all of them positively and Eq. (10) is indeed prepared. [^2] To us, the description of how this kind of sample is obtained in [19] is not transparent. In the following subsections, two classical algorithms for LWE problem is developed assuming the resized sample is given.
5.2 Search on the Error
Given a resized sample, a naive classical strategy is to exhaustively search $s_{j}$ with $O(qn)$ complexity (in which the scenario is indeed discussed in [19]), but more sophisticated search would be to examine the error.
The (measured) resized sample of the $j$th component reads $b_{j}^{\prime}=a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$, $|e_{j}^{\prime}|\leq k^{\prime}=O(n^{3}k)$. Let us consider the following procedure:
  1. 1.
    Set a value of $e_{j}^{\prime}$ such that $|e_{j}^{\prime}|\leq k^{\prime}$.
  2. 2.
    For a measured sample, compute $\tilde{s}_{j}:=(a_{j}^{\prime})^{-1}(b_{j}^{\prime}-e_{j}^{\prime})\bmod q$.
  3. 3.
    For $1\leq i\leq M$ for some integer $M$, test if $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$, where $b_{j,i}^{\prime}$ and $a_{j,i}^{\prime}$ are $i$th sample for the test on the $j$th component.
  4. 4.
    If it passes the test, return $\tilde{s}_{j}$. Otherwise, set a different value for $e_{j}^{\prime}$ and repeat.
Let us estimate the complexity. If $\tilde{s}_{j}=s_{j}$, then $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is always true and it will pass. The probability that a candidate $e_{j}^{\prime}$ leads to the correct $\tilde{s}_{j}$ is thus at least $\Omega(1/k^{\prime})$. On the other hand if $\tilde{s}_{j}\neq s_{j}$, then the probability that $|b_{j,i}^{\prime}-a_{j,i}^{\prime}\tilde{s}_{j}\bmod q|\leq k^{\prime}$ is true is $(2k^{\prime}+1)/q$, thus $\tilde{s}_{j}\neq s_{j}$ will pass with the probability $((2k^{\prime}+1)/q)^{M}$.
Compare it with the algorithm in [19]. In [19], the probability that the BV subroutine outputs $\tilde{s}_{j}=s_{j}$ is $\Omega(1/k^{\prime})$ (Eq. (17) in [19] with $|v_{j}|\approx q$). The probability that $M$-trial test accepts an incorrect answer is at most $((2k^{\prime}+1)/q)^{M}$ as in Eq. (12) in [19]. It is not difficult to see that the above classical procedure is at least as efficient as the algorithm proposed in [19].
5.3 Lattice reduction

### v2

@@ Introduction
We see that the operation is efficiently computable as any power of $\mathbf{P}$ can be classically pre-computed by knowing the fixed polynomial $\phi(X)$.
Now let ${\bm{s}}=\mathsf{V}(s)$ and ${\bm{e}}_{a}=(e_{a,0},\ldots,e_{a,n-1})=\mathsf{V}(e_{a})$. By Lemma 1, Equation (4), and the linearity of $\mathsf{V}$, we have
$\mathsf{V}(s\cdot a+e_{a}\bmod\phi)=\mathsf{V}(s)\cdot\mathsf{M}_{\phi}(a)+\mathsf{V}(e_{a})=\big{(}\!\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0},\left<{\bm{s}},{\bm{a}}_{1}\right>+e_{a,1},\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\big{)}\enspace.$
$\left|a(X)\,\right\rangle\left|0\,\right\rangle\left|s(X)\!\cdot\!a(X)+e_{a}(X)\bmod\phi\,\right\rangle$ $\stackrel{{\scriptstyle\mathsf{V}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{0}}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}_{0}\right>+e_{a,0},\ldots,\left<{\bm{s}},{\bm{a}}_{n-1}\right>+e_{a,n-1}\,\right\rangle$
$\stackrel{{\scriptstyle U_{\phi,j}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{a}}_{j}\,\right\rangle\left|\ldots,\,\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j},\,\ldots\,\right\rangle\enspace,$
where $\bmod\;q$ is omitted in the third register. Uncomputing irrelevant information, we have a quantum sample with secret ${\bm{s}}$:
$\left|{\bm{a}}_{j}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}_{j}\right>+e_{a,j}\bmod q\,\right\rangle\enspace.$
Alternatively, interchanging the role of $a(X)$ and $s(X)$ in the transformation, we have a quantum sample with secret ${\bm{s}}_{j}$:
$\left|a(X)\,\right\rangle\left|0\,\right\rangle\left|a(X)\!\cdot\!s(X)+e_{a}(X)\bmod\phi\,\right\rangle\stackrel{{\scriptstyle\mathsf{V}}}{{\longmapsto}}\left|{\bm{a}}\,\right\rangle\left|{\bm{0}}\,\right\rangle\left|\left<{\bm{a}},{\bm{s}}_{0}\right>+e_{a,0},\ldots,\left<{\bm{a}},{\bm{s}}_{n-1}\right>+e_{a,n-1}\,\right\rangle\enspace,$
where,
${\bm{s}}_{j}=(\mathsf{M}_{\phi}({\bm{s}})_{j})^{T}=\big{(}\!\left<{\bm{s}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{s}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{s}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\enspace.$
In either case, the LWE algorithm is then applied to solve the problem.
Result 2: A classical algorithm for SIS problem
We develop an efficient classical algorithm for solving SIS given that the following quantum sample is provided. In [6, Appendix B], a quantum sample for SIS is defined by,
$\left|\Psi^{\textup{\tiny sis}}\,\right\rangle=\frac{1}{\sqrt{q^{nm}}}\sum_{\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}}\left|\mathbf{A}\,\right\rangle\left|\mathbf{A}{{\bm{v}}}\bmod q\,\right\rangle\enspace,$ (5)
where ${\bm{v}}\in\mathbb{Z}^{n}$ is a short secret vector such that $0<\|{\bm{v}}\|\leq\beta$. The authors have developed a polynomial time quantum algorithm for finding ${\bm{v}}$ by using $O(2n\beta)$ quantum samples. The detailed description of the algorithm is not covered here. Instead, let us focus on the problem itself. From the definition of the SIS problem, we are given an under-determined system of equations ($m<n$). Hardness of the SIS problem comes from the fact that among the exponentially many solutions ($\Omega(q^{n-m})$), there exists no known way to choose a short one within polynomial time.
Assuming the SIS quantum sample in Equation (5) is obtainable, an efficient classical algorithm can be designed. Notice that by measuring a quantum sample, one obtains a classical sample $(\mathbf{A},\mathbf{A}{{\bm{v}}})\in\mathbb{Z}_{q}^{m\times n}\times\mathbb{Z}_{q}^{m}$. Prepare $t=\Omega(n/m)$ classical samples of the form $({\bm{A}}^{(i)},{\bm{z}}^{(i)})\in\mathbb{Z}_{q}^{m\times n}\times\mathbb{Z}_{q}^{m}$ which satisfies ${\bm{A}}^{(i)}{\bm{v}}\equiv{\bm{z}}^{(i)}\bmod q$ for some nonzero vector ${\bm{v}}\in\mathbb{Z}^{n}$ with $\|{\bm{v}}\|\leq\beta$. Each classical sample leads to $m$ linear equations (be reminded that one quantum sample leads to one classical sample, and one classical sample here corresponds to $m$ linear equations),
$A^{(i)}_{11}v_{1}+\cdots+A^{(i)}_{1n}v_{n}\equiv z^{(i)}_{1}\bmod q,\hskip 2.84526pt\ldots\hskip 2.84526pt,\,A^{(i)}_{m1}v_{1}+\cdots+A^{(i)}_{mn}v_{n}\equiv z^{(i)}_{m}\bmod q,$
where $A^{(i)}_{kl}$ is the $k$th row, $l$th column element of $\mathbf{A}^{(i)}$, and $v_{k}$, $z_{k}^{(i)}$ are the $k$th elements of ${\bm{v}}$ and ${\bm{z}}^{(i)}$, respectively. Therefore the total number of linear equations reads $mt=\Omega(n)$. Since there are $n$ unknowns while we have $\Omega(n)$ equations, we can find the secret vector ${\bm{v}}$ by solving the system of equations using Gaussian elimination within $O(n^{3})$ time. The total number of (quantum) samples used is $\Theta(n/m)$.
Result 3: Classical algorithms for size-reduced LWE problem
Song et al. have published a paper in 2022 which improved the result of Grilo et al. assuming a resized sample [8]. In a nutshell, the improved algorithm and Algorithm 2 work similarly except for the resized samples. Therefore, here we drop the details on the algorithm and jump right into the discussion on the resized sample. Recall first that an LWE sample is $({\bm{a}},b=\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$, $|e|\leq k$.
Let $\mathcal{V}_{j}\subseteq\mathbb{Z}_{q}$, ${\bm{s}}=(s_{0},\ldots,s_{n-1})\in\mathbb{Z}_{q}^{n}$, and consider the noise model $\chi_{\textup{\tiny lwe}}$ as a discrete uniform or a bounded Gaussian distribution in the interval $[-\xi,\xi]$ around zero with $\xi\ll q$. Assume a resized quantum sample examined by [8, Equation (9)] is given,
$\big{|}\psi_{\mathcal{V}_{j}}\big{\rangle}=\frac{1}{\sqrt{|\mathcal{V}_{j}|}}\sum_{a_{j}^{\prime}\in\mathcal{V}_{j}}\left|a_{j}^{\prime}\,\right\rangle\left|a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q\,\right\rangle,$ (6)
where $e_{j}^{\prime}$ is derived according to [8, Equation (4)-(8)] satisfying,
$|e_{j}^{\prime}|\leq\xi^{\prime}=O(n^{3}\xi)\ll q\enspace.$
Below two classical algorithms are introduced assuming the resized samples are given.
Search on the error
We first introduce an algorithm that exhaustively searches the ‘correct error’. Given a resized sample, an immediate (classical) strategy would be to exhaustively search $s_{j}$ with $O(qn)$ complexity (in which the scenario is indeed discussed in [8, Discussion]), but a more sophisticated search would be to examine the error.
Consider an oracle $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$ which on query returns the (measured) resized sample of the $j$th component $(a_{j}^{\prime},b_{j}^{\prime})$, where $b_{j}^{\prime}=a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$. Given access to the oracle, Algorithm 4 is developed.
Algorithm 4 $\mathsf{SearchError}$
1:error bound $\xi^{\prime}$, oracle access to $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
2:$\tilde{s}_{j}\in\mathbb{Z}_{q}$
3:$(a_{j}^{\prime},b_{j}^{\prime})\leftarrow\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
4:${\rm e}\leftarrow-\xi^{\prime}$
5:repeat
6:     $\tilde{e}_{j}^{\prime}\leftarrow\rm{e}$
7:     $\tilde{s}_{j}\leftarrow(a_{j}^{\prime})^{-1}(b_{j}^{\prime}-\tilde{e}_{j}^{\prime})\bmod q$
8:     if  $\mathsf{e\textup{-\!}Test}\big{(}\tilde{s}_{j},\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}\big{)}=1$ then
9:         return $\tilde{s}_{j}$
10:     end if
11:     $\rm{e}\leftarrow\rm{e}+1$
12:until ${\rm e}>\xi^{\prime}$
13:return $1$
Algorithm 5 $\mathsf{e\textup{-\!}Test}$
1:$\tilde{s}_{j}\in\mathbb{Z}_{q}$, oracle access to $\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
2:$1$ (pass) or 00 (fail)
3:$i\leftarrow 0$
4:repeat
5:     $(a_{j}^{\prime\prime},b_{j}^{\prime\prime})\leftarrow\mathcal{O}^{\textup{\tiny resized}}_{s_{j},\mathcal{V}_{j}}$
6:     if $|b_{j}^{\prime\prime}-a_{j}^{\prime\prime}\tilde{s}_{j}\bmod q|>\xi^{\prime}$ then
7:         return 00
8:     end if
9:     $i\leftarrow i+1$
10:until $i\geq M$
11:return $1$
If $\tilde{e}_{j}^{\prime}=e_{j}^{\prime}$, then replacing $b_{j}^{\prime}$ in Line 5 by $a_{j}^{\prime}s_{j}+e_{j}^{\prime}\bmod q$ affirms $\tilde{s}_{j}=s_{j}$. Since the error is exhaustively searched within the range $-\xi^{\prime}\leq e_{j}^{\prime}\leq\xi^{\prime}$, Algorithm 4 eventually outputs the correct $\tilde{s}_{j}(=s_{j})$ with time complexity $O(\xi^{\prime})$, as long as the probability that wrong $\tilde{s}_{j}(\neq s_{j})$ passes the test is low enough. If $\tilde{s}_{j}\neq s_{j}$, for a random sample $(a_{j}^{\prime\prime},b_{j}^{\prime\prime})$, the probability that $|b_{j}^{\prime\prime}-a_{j}^{\prime\prime}\tilde{s}_{j}\bmod q|\leq\xi^{\prime}$ holds is $(2\xi^{\prime}+1)/q$. Therefore, such $\tilde{s}_{j}\neq s_{j}$ will pass $\mathsf{e\textrm{-}Test}$ with the probability $((2\xi^{\prime}+1)/q)^{M}$.
In the previous work [8], the probability that the BV subroutine outputs $\tilde{s}_{j}=s_{j}$ is $\Omega(1/\xi^{\prime})$ ([8, Equation (17)] with $|v_{j}|\approx q$), thus $O(\xi^{\prime})$ repetitions of the subroutine would give the correct $\tilde{s}_{j}(=s_{j})$. The probability that their $M$-trial test accepts an incorrect answer is at most $((2\xi^{\prime}+1)/q)^{M}$ as in [8, Equation (12)].
One may notice that the two algorithms have similar upper bounds on the number of repetitions (of error guessing in Algorithm  4 and BV subroutine in  [8, Section 3]). Moreover, the probability that an incorrect candidate passes the $\mathsf{e\textrm{-}Test}$ and $M$-trial test is exactly the same. Direct comparison of the two algorithms is not immediate due to the differences in approaches (e.g., exhaustive or else) and unit computations (quantum or classical), but it is plausible to assess the two as similar in performance. Except for the query, Algorithm 4 works classically.
Lattice reduction
There is another way to find $s_{j}^{\prime}$ using techniques in lattice theory. Given two classical samples $(a_{j,k}^{\prime},b_{j,k}^{\prime}=a_{j,k}^{\prime}s_{j}+e_{j,k}^{\prime}\bmod q)$ for $k=1,2$, consider two-dimensional vectors ${\bm{a}}_{j}^{\prime}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$, ${\bm{b}}_{j}^{\prime}=(b_{j,1}^{\prime},b_{j,2}^{\prime})$, ${\bm{e}}_{j}^{\prime}=(e_{j,1}^{\prime},e_{j,2}^{\prime})$. Then by the construction, we know there is a vector ${\bm{v}}\in\mathbb{Z}^{2}$ such that ${\bm{v}}\equiv s_{j}{\bm{a}}_{j}^{\prime}\bmod q$ and $\|{\bm{b}}_{j}^{\prime}-{\bm{v}}\|=\|{\bm{e}}_{j}^{\prime}\|$. In particular, the vector ${\bm{v}}$ is very close to ${\bm{b}}_{j}^{\prime}$ since $\|{\bm{e}}_{j}^{\prime}\|\leq\sqrt{2}\xi^{\prime}$ is small. Furthermore, notice that a set $\{{\bm{w}}\in\mathbb{Z}^{2}:{\bm{w}}\equiv x{\bm{a}}_{j}^{\prime}\bmod q,x\in\mathbb{Z}\}$ constitutes a lattice in which ${\bm{v}}$ is included. In other words, for given ${\bm{b}}_{j}^{\prime}$, finding $s_{j}$ can be solved by looking for a lattice vector that is closest to ${\bm{b}}_{j}^{\prime}$, known as the closest vector problem (CVP).
In general, the CVP is hard to solve, that is, it takes time exponential in the lattice dimension. However in our case, it is sufficient to solve the CVP over a two-dimensional lattice, which is tractable by using the Babai’s nearest plane algorithm with a short basis computed by the Lagrange-Gauss algorithm.
Lagrange-Gauss algorithm. This algorithm is useful for finding a shortest basis of a lattice in two-dimension. We say that an ordered basis ${\bm{g}}_{1},{\bm{g}}_{2}$ of a lattice $L$ is Lagrange-Gauss reduced if $\|{\bm{g}}_{1}\|\leq\|{\bm{g}}_{2}\|\leq\|{\bm{g}}_{2}+z{\bm{g}}_{1}\|$ for all $z\in\mathbb{Z}$. It can be shown that if a basis ${\bm{g}}_{1},{\bm{g}}_{2}$ of a lattice $L$ is Lagrange-Gauss reduced, then $\|{\bm{g}}_{i}\|=\lambda_{i}(L)$, where $\lambda_{i}(L)$ is the $i$th successive minimum of $L$ [20, Chapter 17]. This basis can be computed efficiently by Algorithm 6.
Algorithm 6 $\mathsf{LagrangeGauss}$
1:basis ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathbb{R}^{2}$ for a lattice $L$ such that $\|{\bm{g}}_{1}\|<\|{\bm{g}}_{2}\|$
2:basis ${\bm{g}}_{1},{\bm{g}}_{2}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(L)$
3:while $\|{\bm{g}}_{1}\|<\|{\bm{g}}_{2}\|$ do
4:     ${\bm{g}}_{2}\leftarrow{\bm{g}}_{2}-\left\lfloor\!\frac{\langle{\bm{g}}_{1},{\bm{g}}_{2}\rangle}{\|{\bm{g}}_{1}\|^{2}}\!\right\rceil{\bm{g}}_{1}$
5:     Swap ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$
6:end while
7:return ${\bm{g}}_{1},{\bm{g}}_{2}$
Detailed analysis of the algorithm can be found in [20, Chapter 17]. The time complexity of the algorithm is known to be $O(\log^{3}(\max_{i}{\|{\bm{g}}_{i}\|^{2}}))$, where ${\bm{g}}_{i}$ are inputs.
Nearest plane algorithm. The algorithm is for solving CVP. Let ${\bm{g}}_{1},\ldots,{\bm{g}}_{n}\in\mathbb{R}^{n}$ be basis vectors of a lattice $L$. Given a target vector ${\bm{t}}\in\mathbb{R}^{n}$, we are asked to find a vector ${\bm{v}}\in L$ such that $\|{\bm{t}}-{\bm{v}}\|$ is minimized.
Babai’s nearest plane algorithm gives an approximate solution to CVP [21] in general dimension. Intuitively, what it does is to fix some basis vectors and then find a linear combination of remaining vectors such that it is closest to the projected ${\bm{u}}$ onto the hyperplane spanned by the fixed basis. Algorithm 7 describes the nearest plane algorithm in dimension $2$. We refer to [20, Chapter 18] for general dimension.
Algorithm 7 $\mathsf{NearestPlane}$
1:basis ${\bm{g}}_{1},{\bm{g}}_{2}$ for a lattice $L\subseteq\mathbb{R}^{2}$, target ${\bm{t}}\in\mathbb{R}^{2}$
2:${\bm{v}}\in L$
3:$\mathbf{g}_{1}^{\ast}\leftarrow{\bm{g}}_{1}$
4:$\mathbf{g}_{2}^{\ast}\leftarrow{\bm{g}}_{2}-\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\mathbf{g}_{1}^{\ast}$
5:${\bm{t}}_{2}\leftarrow{\bm{t}}$
@@
7:${\bm{t}}_{1}\leftarrow{\bm{t}}_{2}-z_{2}{\bm{g}}_{2}$
@@
9:return ${\bm{v}}=z_{1}{\bm{g}}_{1}+z_{2}{\bm{g}}_{2}$
Proposition 1.
Let ${\bm{g}}_{1},{\bm{g}}_{2}\in L$, and ${\bm{t}}\in\mathbb{R}^{2}$ be inputs to Algorithm 7 and let $\mathbf{g}_{1}^{*},\mathbf{g}_{2}^{*}$ be Gram-Schmidt orthogonalized vectors of ${\bm{g}}_{1},{\bm{g}}_{2}$. If there exists ${\bm{w}}\in L$ such that $\|{\bm{t}}-{\bm{w}}\|<\frac{1}{2}\min_{i\in\{1,2\}}(\|\mathbf{g}_{i}^{\ast}\|)$, then Algorithm  7 outputs ${\bm{w}}$.
Proof.
Let ${\bm{v}}=z_{1}{\bm{g}}_{1}+z_{2}{\bm{g}}_{2}$ be the output of Algorithm 7, and let ${\bm{t}}_{1},{\bm{t}}_{2}$ be vectors as in the algorithm. Assume there exists ${\bm{w}}$ such that ${\bm{w}}=w_{1}{\bm{g}}_{1}+w_{2}{\bm{g}}_{2}={\bm{t}}-{\bm{e}}$ for some $w_{1},w_{2}\in\mathbb{Z}$ and $\|{\bm{e}}\|<\frac{1}{2}\min_{i\in\{1,2\}}(\|\mathbf{g}_{i}^{\ast}\|)$. Then we have a useful bound on $|\langle{\bm{e}},\mathbf{g}_{i}^{\ast}\rangle|/\|\mathbf{g}_{i}^{\ast}\|^{2}$ for $i=1,2$,
$\frac{|\langle{\bm{e}},\mathbf{g}_{i}^{\ast}\rangle|}{\|\mathbf{g}_{i}^{\ast}\|^{2}}\leq\frac{\|{\bm{e}}\|\|\mathbf{g}_{i}^{\ast}\|}{\|\mathbf{g}_{i}^{\ast}\|^{2}}<\frac{1}{2}\enspace.$ (7)
Recall that ${\bm{t}}_{2}={\bm{t}}={\bm{w}}+{\bm{e}}$. From
$\frac{\langle{\bm{t}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=\frac{\langle{\bm{w}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=\frac{\langle w_{1}{\bm{g}}_{1}+w_{2}{\bm{g}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}=w_{2}+\frac{\langle{\bm{e}},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}\enspace,$
we see that $z_{2}=\left\lfloor\frac{\langle{\bm{t}}_{2},\mathbf{g}_{2}^{\ast}\rangle}{\|\mathbf{g}_{2}^{\ast}\|^{2}}\right\rceil=w_{2}$ due to Equation (7). Similarly, recalling ${\bm{t}}_{1}={\bm{t}}-z_{2}{\bm{g}}_{2}$ and from
$\frac{\langle{\bm{t}}-z_{2}{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}$ $=\frac{\langle{\bm{w}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}-z_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}$
$=w_{1}+w_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}-z_{2}\cdot\frac{\langle{\bm{g}}_{2},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}=w_{1}+\frac{\langle{\bm{e}},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\enspace,$
we have $z_{1}=\left\lfloor\frac{\langle{\bm{t}}_{1},\mathbf{g}_{1}^{\ast}\rangle}{\|\mathbf{g}_{1}^{\ast}\|^{2}}\right\rceil=w_{1}$, proving ${\bm{v}}={\bm{w}}$. ∎
The time complexity of Algorithm 7 is $O(\log^{2}B)$ where $B=\max\{\|{\bm{g}}_{1}\|,\|{\bm{g}}_{2}\|\}$.
Finding $s_{j}$ by lattice reduction. In addition to the above algorithms, we adopt Gaussian heuristic which predicts that the number of lattice points inside a measurable set $\mathcal{B}\subset\mathbb{R}^{n}$ is approximately equal to $\operatorname{Vol}(\mathcal{B})/\operatorname{Vol}(L)$ where $L$ is a full-rank lattice in $\mathbb{R}^{n}$ and $\operatorname{Vol}(L):=\det(L)$. Gaussian heuristic is widely used in analyzing the performance of lattice-related algorithms. Applied to Euclidean ball in dimension $n$, we expect the length of a nonzero shortest vector in a random lattice $L$ to be approximately,
$\sqrt{\frac{n}{2\pi e}}\det(L)^{1/n}\enspace.$